Find all six trigonometric functions of θ if the point (-15, 36) is on the terminal side of θ. (If an answer is undefined, enter UNDEFINED.)
sin(θ) =

cos(θ) =

tan(θ) =

cot(θ) =

sec(θ) =

csc(θ) =

Answer by Alan3354(69443)

Find all six trigonometric functions of θ if the point (-15,36) is on the terminal side of θ. (If an answer is undefined, enter UNDEFINED.)
----------
There's no need to get excited about UNDEFINED.
---
x = -15
y = 36
r = sqrt(x^2 + y^2) = 39
----
sin = y/r
cos = x/r
tan = y/x
---
cot = 1/tan
sec = 1/cos
csc = 1/sin
==============
Learn this.
Remember this.
Know this.